If s varies directly as T when s=8, t=3 find s when t=12
Found 2 solutions by checkley77, richwmiller:
Answer by checkley77(12844)

8/3=S/12
3S=8*12
3S=96
S=96/3
S=32 ANS.
PROOF:
8/3=32/12
3*32=8*12
96=96

Answer by richwmiller(17219)

t=ks
3=8k
3/8=k
t=ks
12=3/8s
96/3=s
32=s